OpenStudy (anonymous):

how can we find the value of \(\dfrac{1}{-tan135 degree}\)

OpenStudy (anonymous):

@ganeshie8 @mathmath333 @surd @ash2326

OpenStudy (anonymous):

@paki @hartnn

OpenStudy (asnaseer):

what do you know about \(\tan(180-\theta)\) in terms of \(\tan(\theta)\)?

OpenStudy (anonymous):

welcome here after a long time here @asnaseer

OpenStudy (anonymous):

1/-tan135 =?just \(-tan135~dgre\)

OpenStudy (asnaseer):

@ziqbal103 - please answer the question I asked above - its answer will guide you to how to calculate \(\frac{1}{-\tan(135^0)}\)

OpenStudy (anonymous):

yes its answer i want to ask ...as i know the values of 1/tan 2pi ...1/tan 60 degree but don't know its value

OpenStudy (asnaseer):

My question was: What is \(\tan(180-\theta)=\) in terms of \(\tan(\theta)\)?

OpenStudy (asnaseer):

for any \(\theta\)

OpenStudy (anonymous):

for theta =-tan135

OpenStudy (asnaseer):

no. e.g.:\[\sin(180-\theta)=\sin(\theta)\]\[\cos(180-\theta)=-\cos(\theta)\]so:\[\tan(180-\theta)=?\]

OpenStudy (asnaseer):

have you heard of the CAST acronym related to trigonometry?

OpenStudy (anonymous):

what is CAST? sorry i don;t know

OpenStudy (asnaseer):